# A survey of degree-boundedness

### v1

@@ 4.1 Dichotomy
Consider a graph $G$ in $\mathcal{F}$ with biclique number $\tau$ whose average degree exceeds the desired degree bound. If we are aiming for the degree bound of $d$, then we may assume that the graph $G$ is roughly $d$-degenerate, since otherwise we could instead consider a smaller induced subgraph of $G$ whose average degree still exceeds the bound. The proofs of all the three mentioned results aim for finding a $4$-cycle-free induced subgraph of $G$ with average degree at least $k$, which leads to a contradiction as $\mathcal{F}$ is hereditary. Given that $G$ has large average degree, the problem can be separated into the following two cases:
@@
The general approach to prove the above was already used earlier by Kwan, Letzter, Sudakov, and Tran [87]. One of the improvements made in [44] was to use an approach of Montgomery, Pokrovskiy, and Sudakov [101] to “bump” how large $A$ is in comparison to $B$. So there is an intermediate step where a “biregular” graph of large average degree is found; this in turn guarantees the existence of an “almost regular” induced subgraph of large degree; see [78, Lemma 3.7] in a paper of Janzer and Sudakov which resolved the Erdős-Sauer problem.
Since this type of dichotomy is (at least at a broad level) a common element of all of the approaches, we focus instead on the key new improvements of Bourneuf, Bucić, Cook, and Davies [17], and of Girão and Hunter [67] which were used to take care of the two cases. Subsections 4.2 and 4.3 are devoted to introducing the techniques used to handle the two respective cases. Here is a more rigorous statement of the type of dichotomy theorem that is often used. Note that in case (B), the graph is not bipartite yet. This can be dealt with later on using degeneracy.
Theorem 4.3.
[44, Lemma 3.4] Let $L,d\geqslant 16$ be positive numbers. Then for any $n$-vertex graph $G$ with average degree $d$ that is $d$-degenerate, either:
@@ 4.2 Almost regular graphs
In fact, Girão and Hunter [67, Lemma 7.1] proved that when $t$ is sufficiently large in comparison to the number of vertices of $H$, that we can take $\epsilon_{H}=1/(100\cdot\Delta(H))$. The following lemma, also proven by Girão and Hunter, gives a good choice of the graph $H$. For any integers $\ell$ and $r$ with $r\leqslant\ell$, we write $K_{\ell}^{(r)}$ for the bipartite graph with $\ell$ vertices on one side and $\binom{\ell}{r}$ vertices on the other side, where for each $r$-subset of the $\ell$ vertices, there exists a vertex whose neighborhood is precisely that $r$-set. (An $r$-set is just a set of size $r$.)
Lemma 4.4 ([67, Lemma 7.4]).
@@
We may assume that $k\geqslant 2$. So let us now fix a graph $H$ as above. Let $\epsilon_{H}$ and $p_{H}(t)$ be as in Theorem 2.11. In fact, as mentioned before, we can take $\epsilon_{H}=1/(100\cdot\Delta(H))\geqslant 1/(200k)$. So $\epsilon_{H}$ is even relatively large. Note that $H$ is not an induced subgraph of any graph in $\mathcal{F}$, and so we may apply Theorem 2.11 to such graphs. We suppose that case (A) of the dichotomy theorem (Theorem 4.3) holds. Thus, $G_{0}$ is a graph in $\mathcal{F}$ which has average degree $d_{0}$ at least $\Omega(d/\log^{2}(L))$ and maximum degree $\Delta_{0}$ at most $O(d_{0}\cdot\log^{2}(L))$. We let $t$ be the smallest integer so that $G_{0}$ has no induced subgraph isomorphic to $K_{t,t}$. When this dichotomy theorem is applied, we will have that $\log^{2}(L)$ is still fairly small in comparison to $d$. So the maximum degree will be not too much larger than the average degree.
The proof proceeds by finding an induced subgraph of $G_{0}$ by including vertices independently at random with some probability $q$ and then deleting the vertices in any $4$-cycle. The resulting graph will be $4$-cycle-free and will, in expectation, have average degree more than $k$. In this manner we obtain a contradiction. The key observation is that Theorem 2.11 can be used to obtain a rather efficient bound on the number of $4$-cycles. The point is that for every edge $xy\in E(G_{0})$, by applying Theorem 2.11 to the subgraph of $G_{0}$ induced on the neighbors of $x$ and $y$, we know there are at most $8p_{H}(t)\Delta_{0}^{2-\epsilon_{H}}$-many 4-cycles containing the edge $xy$ in $G_{0}$. This is the main idea and turns out to be enough for the proof to go through.
@@ 4.3 Unbalanced bipartite graphs
An argument using the $d$-degeneracy of $G$ shows that at least $|A|/20$ vertices in $A$ have at least $d/20$ and at most $10d$ neighbors in $B$; among those vertices, since $G$ is $d$-degenerate, there exists an independent subset $A_{0}$ of size $|A_{0}|\geqslant(|A|/20)/(d+1)\geqslant|A|/40d$. We then pass to the induced subgraph $G_{0}$ of $G$ which has vertex-set $A_{0}\cup B$, where $|A_{0}|>\frac{L}{40d}|B|$, $A_{0}$ is independent in $G_{0}$, and every vertex in $A_{0}$ has degree between $d/20$ and $10d$ in the graph $G_{0}$.
@@
Before continuing with the proof sketch of the polynomial bound, we briefly mention the different technique used in the proof of the singly exponential bound (Theorem 4.1), and reveal the choice of $L$. The proof of Theorem 4.1 applies the dichotomy with $L=2^{d^{\delta}}$, where $\delta=1/200t$. A probabilistic argument similar to the one in Subsection 4.2 resolves Case (A) by applying the classical extremal bound from the Kővári-Sós-Turán Theorem[44, Lemma 3.3]. For Case (B), the proof is done by first passing to an induced bipartite graph $G[A^{\prime}\cup B^{\prime}]$ where $A^{\prime}\subset A,B^{\prime}\subset B$ are independent and every vertex in $A^{\prime}$ has the same degree, and then applying Füredi’s result to find an induced $4$-cycle-free subgraph with average degree at least $k$ in it[44, Theorem 5.1].
@@
In the proof of the Theorem 4.2, the dichotomy is performed with $L:=d^{D+3}$, where $D$ is an integer such that every $K_{k,k}$-free graph in $\mathcal{F}$ has average degree less than $D$, which exists because $\mathcal{F}$ is degree bounded. The proof sketch in Subsection 4.2 for case (A) goes through as $d_{0}=\Omega(d/\log^{2}(L))\geqslant{O}(d^{1/2})$ and $\Delta/d_{0}=O((\log^{2}(L))=O((D\log(d))^{2})\leqslant{O}(d_{0}^{1/1000k})$, when $d$ is large enough.
@@
We say that a collection $\mathcal{S}$ of subsets of some ground set, in our case the set of vertices $B$, $k$-shatters a set $R\subseteq B$ if for every $k$-subset $S\subseteq R$, there exists a set in $\mathcal{S}$ whose intersection with $R$ is precisely $S$. Given the graph $G_{0}$, consider a vertex $x\in A_{0}$. For any independent subset $J$ contained in the neighborhood of $x$, if the collection $\{N(x^{\prime})\cap N(x):x\in A_{0}\setminus\{x\}\}$ $2k$-shatters a $8k^{2}$-subset of $J$, then $G_{0}$ contains the graph $K_{8k^{2}}^{(2k)}$ as desired.
@@
The proof utilizes the following “shattering lemma". The idea is to apply this lemma within independent sets contained in the neighborhood of vertices $x$. If the lemma does not apply, then instead we find an induced subgraph where certain pairs of vertices have few common neighbors. This will also let us remove all big bicliques and find another win.
Lemma 4.5.
@@
[67, Lemma 8.4] Let $D,t\geqslant 2$ and $h$ be integers. Then there exists $\delta>0$ such that the following holds when $n$ is large enough. For any collection $\mathcal{S}$ of subsets of $\{1,2,\ldots,n\}$ with $F<\delta n$ for all $F\in\mathcal{S}$, if $\mathcal{S}$ does not $t$-shatter an $h$-subset of $\{1,2,\ldots,n\}$, then there exists $I\subseteq\{1,2,\ldots,n\}$ with $|I|=D$ such that $|I\cap S|<k$ for all $S\in\mathcal{S}$.
@@
By arguments similar to case (A), one can prove that we can find large independent sets in the neighborhood of each vertex. The proof is finished by using a probabilistic argument to reduce from having relatively few common neighbors among independent sets within the neighborhoods of vertices to actually being $K_{k,k}$-free while still having average degree at least $D$. This broad overview outlines some of the improvements used to obtain a polynomial degree-bounding function.
@@ 5 Examples
We note that every graph $G$ is an intersection graph of a collection of subsets of $V(G)$. For each vertex $v$ of $G$, let $B_{v}\subseteq V(G)$ be the set containing $v$ and all of its neighbors. Then two vertices $u$ and $v$ are adjacent in $G$ if and only if $B_{u}$ and $B_{v}$ intersect. (We say that two sets intersect if their intersection is non-empty.) So in order to obtain a class of graphs that is degree-bounded, we need to impose some structure on how sets in $\mathcal{B}$ can intersect. We discuss a number of ways to use geometry to impose this structure.
@@ 5.1 Convex sets in Euclidean space
We also refer the reader to the proof of Brightwell and Scheinerman [20], which is stated in accessible language and finds a similtaneous representation of the planar dual.
@@
Thus the following lemma is connected both to degree-boundedness and to the minimum degree of planar graphs. In particular, the lemma implies that for any fixed $d$, the class of intersection graphs of balls in $\mathbb{R}^{d}$ is degree-bounded by the linear function $f_{d}(\tau)=(2\tau+1)2^{d}$. This lemma is well-known, but since the proof is short, we provide it anyways.
@@
For any positive integers $k$ and $d$ and any finite $k$-thin collection $\mathcal{B}$ of balls in $\mathbb{R}^{d}$, the intersection graph $G(\mathcal{B})$ has a vertex of degree less than $k\cdot 2^{d}$.
@@
Let $B\in\mathcal{B}$ be a ball of minimum volume; we will show that $B$ is the desired vertex of small degree. By translating and then scaling, we may assume that $B$ is the unit ball centered at the origin, that is, $B=\{x\in\mathbb{R}^{d}:\lVert x\rVert_{2}\leqslant 1\}$. (A unit ball is any ball of radius $1$.) Now note that for any $B^{\prime}\in\mathcal{B}\setminus\{B\}$ which intersects $B$, there exists a unit ball which is contained in $B^{\prime}$ and still intersects $B$. Thus, there exists a $k$-thin collection $\mathcal{B}^{\prime}$ of unit balls so that 1) the degree of $B$ in the intersection graph $G(\mathcal{B})$ is less than $|\mathcal{B}^{\prime}|$, and 2) every ball in $\mathcal{B}^{\prime}$ is a subset of the ball of radius $2$ centered at the origin.
Let us write $v_{d}$ for the volume of a unit ball in $\mathbb{R}^{d}$. It follows from the definition of $\mathcal{B}^{\prime}$ that the sum of the volumes of the balls in $\mathcal{B}^{\prime}$ is $v_{d}|\mathcal{B}^{\prime}|$, and that this sum is at most $k(v_{d}\cdot 2^{d})$ since $v_{d}\cdot 2^{d}$ is the volume of a ball of radius $2$. So $|\mathcal{B}^{\prime}|\leqslant k\cdot 2^{d}$, and the lemma follows. ∎
It is an interesting but likely very difficult question to find the optimal constant which can replace $2^{d}$ in Lemma 5.1. While there are many ways to formulate this problem, let us pose the following version.
@@
While we do not know of any direct reduction, this problem seems highly related to the famous question about the density of sphere packings; see [30].
@@ 5.2 Circle graphs and vertex-minors
The class of circle graphs is closed under taking vertex-minors, and Bouchet [16] found the forbidden vertex-minors: there are only three of them. Surprisingly, this theorem is related to the famous theorem of Kuratowski and Menger about the forbidden minors for planar graphs. The connection was discovered by de Fraysseix [40], and Geelen and Oum [64] have proven a common generalization of the two theorems. Motivated by a structural question of Geelen about classes of graphs which forbid a vertex-minor (see the second author’s thesis for a formal statement of the conjecture [97]), we conjecture the following.

### v2

@@ 4.1 Dichotomy
Aiming for a contradiction, suppose that there is a graph $G$ in $\mathcal{F}$ with biclique number $\tau$ whose average degree exceeds the desired degree bound. If we are aiming for the degree bound of $d$, then we may assume that the graph $G$ is roughly $d$-degenerate, since otherwise we could instead consider a smaller induced subgraph of $G$ whose average degree still exceeds the bound. The proofs of all three mentioned results find a $4$-cycle-free induced subgraph of $G$ with average degree at least $k$, which leads to a contradiction as $\mathcal{F}$ is hereditary. Given that $G$ has large average degree, the problem can be separated into the following two cases:
@@
A similar approach was already used earlier by Kwan, Letzter, Sudakov, and Tran [93]. One of the improvements made in [46] was to use an approach of Montgomery, Pokrovskiy, and Sudakov [107] to “bump” how large $A$ is in comparison to $B$. So there is an intermediate step where a “biregular” graph of large average degree is found; this in turn guarantees the existence of an “almost regular” induced subgraph of large degree; see [83, Lemma 3.7] in a paper of Janzer and Sudakov which resolved the Erdős-Sauer problem.
Since this type of dichotomy is (at least at a broad level) a common element of all of the approaches, we focus instead on the key new improvements of Bourneuf, Bucić, Cook, and Davies [19], and of Girão and Hunter [72] which were used to take care of the two cases. Subsections 4.2 and 4.3 are devoted to introducing the techniques used to handle the two respective cases. Here is a more rigorous statement of the type of dichotomy theorem that is often used. Note that in case (B) of the theorem below, the graph is not bipartite yet. This can be dealt with later on using degeneracy.
Theorem 4.3 ([72, Lemma 3.4]).
Let $L,d\geqslant 16$ be positive integers. Then for any $n$-vertex graph $G$ with average degree $d$ that is $d$-degenerate, either:
@@ 4.2 Almost regular graphs
In fact, Girão and Hunter [72, Lemma 7.1] proved that when $t$ is sufficiently large in comparison to the number of vertices of $H$, we can take $\epsilon_{H}=1/(100\cdot\Delta(H))$. The following lemma, also proven by Girão and Hunter, gives a good choice of the graph $H$. For any integers $\ell$ and $r$ with $r\leqslant\ell$, we write $K_{\ell}^{(r)}$ for the bipartite graph with $\ell$ vertices on one side and $\binom{\ell}{r}$ vertices on the other side, where for each $r$-subset of the $\ell$ vertices, there exists a vertex whose neighborhood is precisely that $r$-set. (An $r$-set is just a set of size $r$.)
Lemma 4.4 ([72, Lemma 7.4]).
@@
We may assume that $k\geqslant 2$. So let us now fix a graph $H$ as above. Let $\epsilon_{H}$ and $p_{H}(t)$ be as in Theorem 2.11. In fact, as mentioned before, we can take $\epsilon_{H}=1/(100\cdot\Delta(H))\geqslant 1/(200k)$. So $\epsilon_{H}$ is even relatively large. Note that $H$ is not an induced subgraph of any graph in $\mathcal{F}$, and so we may apply Theorem 2.11 to such graphs. We suppose that case (A) of the dichotomy theorem (Theorem 4.3) holds. Thus, $G_{0}$ is a graph in $\mathcal{F}$ which has average degree $d_{0}$ at least $\Omega(d/\log^{2}(L))$ and maximum degree $\Delta_{0}$ at most $O(d_{0}\cdot\log^{2}(L))$. We let $t$ be the smallest integer so that $G_{0}$ has no subgraph isomorphic to $K_{t,t}$. When this dichotomy theorem is applied, we will have that $\log^{2}(L)$ is still fairly small in comparison to $d$. So the maximum degree will not be too much larger than the average degree.
The proof proceeds by finding an induced subgraph of $G_{0}$ by including vertices independently at random with some probability $q$ and then deleting the vertices in any $4$-cycle. The resulting graph will be $4$-cycle-free and will, in expectation, have average degree more than $k$. In this manner we obtain a contradiction. The key observation is that Theorem 2.11 can be used to obtain a rather efficient bound on the number of $4$-cycles. The point is that for every edge $xy\in E(G_{0})$, by applying Theorem 2.11 to the subgraph of $G_{0}$ induced on the neighbors of $x$ and $y$, we know there are at most $8p_{H}(t)\Delta_{0}^{2-\epsilon_{H}}$-many 4-cycles containing the edge $xy$ in $G_{0}$. (This induced subgraph has at most $2\Delta_{0}$-many vertices, and each of its edges contributes at most two such $4$-cycles.) This is the main idea and turns out to be enough for the proof to go through.
@@ 4.3 Unbalanced bipartite graphs
An argument using the $d$-degeneracy of $G$ shows that at least $|A|/20$ vertices in $A$ have at least $d/20$ and at most $10d$ neighbors in $B$; among those vertices, since $G$ is $d$-degenerate, there exists an independent subset $A_{0}$ of size $|A_{0}|\geqslant(|A|/20)/(d+1)\geqslant|A|/40d$. We then pass to the induced subgraph $G_{0}$ of $G$ which has vertex-set $A_{0}\cup B$, where $|A_{0}|\geqslant\frac{L}{40d}|B|$, $A_{0}$ is independent in $G_{0}$, and every vertex in $A_{0}$ has degree between $d/20$ and $10d$ in the graph $G_{0}$.
@@
Before continuing with the proof sketch of the polynomial bound, we briefly mention the different technique used in the proof of the singly exponential bound (Theorem 4.1), and reveal the choice of $L$. The proof of Theorem 4.1 applies the dichotomy with $L=2^{d^{\delta}}$, where $\delta=1/200t$ and $t$ is the smallest integer such that $G$ has no subgraph isomorphic to $K_{t,t}$. A probabilistic argument similar to the one in Subsection 4.2 resolves case (A) by applying the classical extremal bound from the Kővári-Sós-Turán Theorem [46, Lemma 3.3]. For case (B), the proof is done by first passing to an induced bipartite subgraph $G[A^{\prime}\cup B^{\prime}]$ where $A^{\prime}\subseteq A,B^{\prime}\subseteq B$ are independent and every vertex in $A^{\prime}$ has the same degree. Then the proof proceeds by applying a theorem of Füredi [63] about uniform hypergraphs to find an induced $4$-cycle-free subgraph with average degree at least $k$ [46, Theorem 5.1].
@@
In the proof of Theorem 4.2, the dichotomy is instead performed with $L:=d^{D+3}$, where $D$ is an integer such that every $K_{k,k}$-free graph in $\mathcal{F}$ has average degree less than $D$, which exists because $\mathcal{F}$ is degree-bounded. The proof sketch in Subsection 4.2 for case (A) goes through as $d_{0}=\Omega(d/\log^{2}(L))\geqslant\Omega(d/(D\log{d})^{2})$ and $\Delta/d_{0}=O(\log^{2}(L))=O((D\log(d))^{2})\leqslant{O}(d_{0}^{1/1000k})$, when $d$ is large enough.
@@
We say that a collection $\mathcal{S}$ of subsets of some ground set, in our case the set of vertices $B$, $k$-shatters a set $R\subseteq B$ if for every $k$-subset $S\subseteq R$, there exists a set in $\mathcal{S}$ whose intersection with $R$ is precisely $S$. Given the graph $G_{0}$, consider a vertex $x\in A_{0}$. For any independent subset $J$ contained in the neighborhood of $x$, if the collection $\{N(x^{\prime}):x^{\prime}\in A_{0}\setminus\{x\}\}$ $2k$-shatters an $8k^{2}$-subset of $J$, then $G_{0}$ contains the graph $K_{8k^{2}}^{(2k)}$ as desired.
@@
The rest of the proof utilizes the following “shattering lemma". The idea is to apply this lemma within independent sets contained in the neighborhood of vertices $x\in A_{0}$. If the lemma does not apply, then instead we find an induced subgraph where certain pairs of vertices have few common neighbors. This will also let us remove all big bicliques and find another win.
Lemma 4.5 ([72, Lemma 8.4]).
@@
Let $D,t\geqslant 2$ and $h$ be integers. Then there exists $\delta>0$ such that the following holds when $n$ is large enough. For any collection $\mathcal{S}$ of subsets of $\{1,2,\ldots,n\}$ with $|F|<\delta n$ for all $F\in\mathcal{S}$, if $\mathcal{S}$ does not $t$-shatter an $h$-subset of $\{1,2,\ldots,n\}$, then there exists $I\subseteq\{1,2,\ldots,n\}$ of size $D$ such that $|I\cap F|<t$ for all $F\in\mathcal{S}$.
@@
By arguments similar to case (A), one can prove that we can find large independent sets in the neighborhood of each vertex. The proof is finished by using a probabilistic argument to find an induced subgraph which is $K_{k,k}$-free yet still has average degree at least $D$. This completes the broad overview of some techniques used to obtain a polynomial degree-bounding function.
@@ 5 Examples
We note that every graph $G$ is an intersection graph of a collection of subsets of $E(G)$. For each vertex $v$ of $G$, let $B_{v}\subseteq E(G)$ be the set of all edges of $G$ that are incident to $v$. Then two vertices $u$ and $v$ are adjacent in $G$ if and only if $B_{u}$ and $B_{v}$ intersect. (We say that two sets intersect if their intersection is non-empty.) So in order to obtain a class of graphs that is degree-bounded, we need to impose some structure on how the sets in $\mathcal{B}$ can intersect. We discuss a number of ways to use geometry to impose this structure.
@@ 5.1 Convex sets in Euclidean space
We also refer the reader to the proof of Brightwell and Scheinerman [22], which is stated in accessible language and finds a simultaneous representation of the planar dual.
@@
Thus the following lemma is connected both to degree-boundedness and to the minimum degree of planar graphs. In particular, the lemma implies that for any fixed $d$, the class of intersection graphs of balls in $\mathbb{R}^{d}$ is degree-bounded by the linear function $f_{d}(\tau)=(2\tau+1)3^{d}$. This lemma is well-known, but since the proof is short, we provide it anyways.
@@
For any positive integers $k$ and $d$ and any finite $k$-thin collection $\mathcal{B}$ of balls in $\mathbb{R}^{d}$, the intersection graph $G(\mathcal{B})$ has a vertex of degree less than $k\cdot 3^{d}$.
@@
Let $B\in\mathcal{B}$ be a ball of minimum volume; we will show that $B$ is the desired vertex of small degree. By translating and then scaling, we may assume that $B$ is the unit ball centered at the origin, that is, $B=\{x\in\mathbb{R}^{d}:\lVert x\rVert_{2}\leqslant 1\}$. (A unit ball is any ball of radius $1$.) Now note that for any $B^{\prime}\in\mathcal{B}\setminus\{B\}$ which intersects $B$, there exists a unit ball which is contained in $B^{\prime}$ and still intersects $B$. Thus, there exists a $k$-thin collection $\mathcal{B}^{\prime}$ of unit balls so that 1) the degree of $B$ in the intersection graph $G(\mathcal{B})$ is less than $|\mathcal{B}^{\prime}|$, and 2) every ball in $\mathcal{B}^{\prime}$ is a subset of the ball of radius $3$ centered at the origin.
Let us write $v_{d}$ for the volume of a unit ball in $\mathbb{R}^{d}$. It follows from the definition of $\mathcal{B}^{\prime}$ that the sum of the volumes of the balls in $\mathcal{B}^{\prime}$ is $v_{d}|\mathcal{B}^{\prime}|$, and that this sum is at most $k(v_{d}\cdot 3^{d})$ since $v_{d}\cdot 3^{d}$ is the volume of a ball of radius $3$. So $|\mathcal{B}^{\prime}|\leqslant k\cdot 3^{d}$, and the lemma follows. ∎
It is an interesting but likely very difficult question to find the optimal constant which can replace $3^{d}$ in Lemma 5.1. While there are many ways to formulate this problem, let us pose the following version.
@@
While we do not know of any direct reduction, this problem seems highly related to the famous question about the density of sphere packings; see [32]. We suspect that $c_{d}$ must be exponential in $d$; perhaps this can be shown by including points of a very fine grid independently at random with an appropriately chosen probability.
@@ 5.2 Circle graphs and vertex-minors
The class of circle graphs is closed under taking vertex-minors, and Bouchet [18] found the forbidden vertex-minors: there are only three of them. Surprisingly, this theorem is related to the famous theorem of Kuratowski and Wagner about the forbidden minors for planar graphs. The connection was discovered by de Fraysseix [42], and Geelen and Oum [69] have proven a common generalization of the two theorems. Motivated by a structural question of Geelen about classes of graphs which forbid a vertex-minor (see the second author’s thesis for a formal statement of the conjecture [103]), we conjecture the following.